OpenStudy (anonymous):

How to show that the standard normal distribution N(0, 1) has CF e^-t(sqr)/2.

OpenStudy (amistre64):

youd have to use the general CF function itslef and sub in those parts

OpenStudy (amistre64):

\[\int _{-inf}^{inf}\frac{e^{something}}{2\sqrt{pi}}\]

OpenStudy (amistre64):

http://mathpages.com/home/kmath045/kmath045.htm

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

yep